# RAG-QA Arena: Evaluating Domain Robustness for Long-form Retrieval Augmented Question Answering

### v1

@@ Abstract
Question answering based on retrieval augmented generation (RAG-QA) is an important research topic in NLP and has a wide range of real-world applications. However, most existing datasets for this task are either constructed using a single source corpus or consist of short extractive answers, which fall short of evaluating large language model (LLM) based RAG-QA systems on cross-domain generalization. To address these limitations, we create Long-form RobustQA (Lfrqa), a new dataset comprising human-written long-form answers that integrate short extractive answers from multiple documents into a single, coherent narrative, covering 26K queries and large corpora across seven different domains. We further propose Rag-qa Arena by directly comparing model-generated answers against Lfrqa’s answers using LLMs as evaluators. We show via extensive experiments that Rag-qa Arena and human judgments on answer quality are highly correlated. Moreover, only 41.3% of the most competitive LLM’s answers are preferred to Lfrqa’s answers, demonstrating Rag-qa Arena as a challenging evaluation platform for future research. [^1] Code: https://github.com/awslabs/rag-qa-arena
RAG-QA Arena: Evaluating Domain Robustness for Long-form Retrieval Augmented Question Answering
@@ 1 Introduction
Traditional reading comprehension task is constrained to fixed contexts Rajpurkar et al. (2016); Kočiský et al. (2018); Huang et al. (2019). It is inadequate at addressing real-world questions, with no context provided for a system to find answers. Such open-ended questions require a system to identify answers in an enormous knowledge base (e.g., Wikipedia) that is computationally prohibitive to feed into LLMs. Retrieval-augmented generative question answering (RAG-QA) becomes an effective tool to filter out massive amounts of noise and select only a few highly relevant passages for LLM-based QA models.
The wide applications of RAG-QA Gao et al. (2023) necessitate the evaluation of systems’ out-of-domain (OOD) performances because a real-world system often confronts new data unseen during training. Existing popular benchmark datasets such as Natural Questions (NQ) Kwiatkowski et al. (2019) and TriviaQA Joshi et al. (2017) consist solely of Wikipedia or Web documents, which fall short of measuring OOD performances.
Robustqa Han et al. (2023) was the first dataset created to benchmark cross-domain robustness for RAG-QA. However, as illustrated by the yellow highlights in Figure 1, Robustqa follows NQ’s annotation format with short answer spans extracted from the documents. Such data format is not the most suitable reference answer to evaluate the current leading LLMs that typically generate long-form responses with multiple information combined in one coherent narrative Brown et al. (2020); Ouyang et al. (2022); OpenAI (2024). Consequently, token overlap metrics used in the extractive QA era Karpukhin et al. (2020); Izacard et al. (2024) will penalize unfairly on the additional supporting tokens generated by LLMs, resulting in extremely low overlap scores. As an example, Fig. 1’s extractive answers have poor Exact-Match or $F_{1}$ scores with the final long-form answer. To create a long reference answer, one can simply concatenate these short answers, but the synthesized answers are either incoherent or ill-formatted, as illustrated by examples in Sec. 3.1.
To address the drawbacks, we propose long-form RobustQA (Lfrqa) that integrates multiple short extractive answers into a coherent long-form answer. Figure 1 shows an annotation where three extractive answers are combined by annotators to create a comprehensive answer. Table 1 summarizes seven features in Lfrqa  that make it uniquely beneficial for RAG-QA evaluations. ASQA Stelmakh et al. (2022) and ELI5 Fan et al. (2019) are the most similar datasets to Lfrqa. However, they are either not directly annotated against the underlying corpus (thus, not RAG-QA), or rely on single-domain data, which is insufficient to benchmark systems’ cross-domain performances.
@@ 5 Experimental Setup
Table 4: Pairwise comparisons between Robustqa, Lfrqa, and LLMs. GPT-4: GPT-4-0125-preview; Mixtral: Mixtral-8x22b-Instruct; Llama-3: Llama-3-70b-Instruct. Answer generated based on top 5 passages using ColBERT-v2. All numbers are % except for Pearson Corr. and Cohen’s Kappa. *** indicates strong correlation with p-values $\ll 0.001$.
@@ 6 Results and Analysis
6.1 LFRQA v.s. RobustQA
[…]
6.2 Quality of Model-based Evaluator
6.3 RAG-QA Arena
Dashboard leaders.  GPT-4o leads the dashboard, with GPT-4-turbo and Mixtral-8x22b-Instruct as close runner-ups. GPT-4o performs the best for [BI], [LI], [RE] and [WR] domains, Mixtral-8x22b-Instruct leads in [FI] and [SC], and GPT-4-turbo champions in [TE].
Note that these results raise a research question about the impact of prompt engineering. We emphasize that the goal of Rag-qa Arena is to propose a reliable evaluation framework, not to conduct extensive prompt engineering or model training to pursue the best RAG-QA system. We provide a dev set of Lfrqa, which can be leveraged in future research for model development purposes.
Elo rating.  Table 3 shows dashboard results of win and win+tie ratio against Lfrqa. We can further convert these pairwise comparisons into Elo ratings similar to Chatbot Arena Chiang et al. (2024). Appendix Table 6 reports Elo ranking. Column (A) uses the same data in Table 3; Column (B) and (C) add pairwise comparisons for all unique model pairs on 700 and 1400 randomly sampled queries across seven domains. So the total added preference pairs are $\frac{N\times K\times(K-1)}{2}$, where K is the number of models in Table 3, and N=700 and 1400 for Column (B) and (C), respectively. More details of Elo ratings and their confidence intervals (CI) are shown in Appendix A.1.
We rank these pairs with the same LLM evaluator and found that in (C), Elo ranking’s 95% CI can separate different LLMs. With a 50% increase of total pairs (and thus the compute), it differs only slightly with the win+tie ranking in Table 3 by flipping the order of llama-3-70b and Command R+. These results present additional evidence that our approach of using Lfrqa only for pairwise comparisons is both efficient and reliable.
Impact of model sizes.  For the non-GPT LLM family, more parameters lead to better performances, but a larger increase in sizes does not always indicate greater performance gains. For example, the two Qwen models have the second-largest difference, but the lift from the smaller to the larger model is marginal. We leave rigorous investigations to future research.
6.4 LLM as Annotators
Citation
Completeness Accuracy Helpfulness
Lfrqa 90.8 88.9 48.1
GPT-4 75.3 65.5 35.2
Using large language models to provide annotations has been explored in previous works Tan et al. (2024). It could provide a more scalable solution than human annotations but can suffer from hallucination and accuracy issues that require human validations Huang et al. (2023). We also experiment with LLM as annotators before we start human annotations. We subsample 100 queries from Lfrqa and prompt GPT-4-0125-preview to follow the similar procedure in Sec. 3.1 to combine answers (Appendix Table 12). Then we request our data linguists to compare Lfrqa and GPT-4 annotations based on 1) Completeness: whether all Robustqa answers are integrated into the final answers; 2) Citation Accuracy: whether citations in answers pointing to the right documents; 3) Helpfulness: defined the same as in Sec. 4. Table 5 shows Lfrqa out-performs GPT-4 annotations by 15.5%, 23.4% and 12.9% for the three dimensions, respectively, suggesting human annotations are both valuable and necessary for our task.
7 Related Work
Longform QA datasets  have been proposed in prior work. ELI5 Fan et al. (2019) and LongFact Wei et al. (2024) contain answers that are either not annotated directly on the corpus, and or not created by humans. Krishna et al. (2021) also points out that ELI5’s small validation set has significant leakage from its train set. ASQA Stelmakh et al. (2022) is the most similar data to our work, but its corpus is in the single Wikipedia domain. Lfrqa is by far the only RAG-QA dataset with comprehensive long-form answers.
8 Conclusion
Limitations
Ethics Statement
Appendix A Appendix
@@
Rating 95% CI Votes Rating 95% CI Votes Rating 95% CI Votes
@@
A.1 ELO Rating
Table 6 shows the results by converting pairwise comparisons to Elo rating Chiang et al. (2024) with Bradley-Terry Model Bradley and Terry (1952). 95% confidence interval (CI) is estimated using bootstrapping. Below, we explain each column in more detail.
(A)  Lfrqa only pairs: we use the same data as in Table 3. All pairs include an answer from Lfrqa.
(B)  Lfrqa + 700 complete pairs: data in (A) + comparisons of all unique pairs of models for 700 randomly sample queries (100 for each domain).
(C)  Lfrqa  + 1400 complete pairs: similar to (B), but increase query number from 700 to 1400.
We can see in (C), that 95% CI can separate different models, so we stop adding more paired comparisons. Notice that this Elo ranking aligns well with ranking results using win+tie in Table 3.
A.2 Dev Set of LFRQA
@@ A.4 Annotation Interface
A.5 Answer Generation Prompts
@@ A.6 Evaluation Interface
A.7 Human Evaluation Instructions
@@
A.8 Evaluation Model Selection
@@ A.9 No Answer Ratio
A.10 Pairwise Evaluation Prompts

### v2

# RAG-QA Arena: Evaluating Domain Robustness for Long-Form Retrieval-Augmented Question Answering

@@ Abstract
Question answering based on retrieval-augmented generation (RAG-QA) is an important research topic in NLP and has a wide range of real-world applications. However, most existing datasets for this task are either constructed using a single source corpus or consist of short extractive answers, which fall short of evaluating large language model (LLM) based RAG-QA systems on cross-domain generalization. To address these limitations, we create Long-form RobustQA (Lfrqa), a new dataset comprising human-written long-form answers that integrate short extractive answers from multiple documents into a single, coherent narrative, covering 26K queries and large corpora across seven different domains. We further propose Rag-qa Arena by directly comparing model-generated answers against Lfrqa’s answers using LLMs as evaluators. We show via extensive experiments that Rag-qa Arena evaluation and human judgments on answer quality are highly correlated. Moreover, only 41.3% of the most competitive LLM’s answers are preferred to Lfrqa’s answers, demonstrating Rag-qa Arena as a challenging evaluation platform for future research. [^1] Code: https://github.com/awslabs/rag-qa-arena
RAG-QA Arena: Evaluating Domain Robustness for Long-Form Retrieval-Augmented Question Answering
@@ 1 Introduction
Traditional reading comprehension task is constrained to fixed contexts Rajpurkar et al. (2016); Kočiský et al. (2018); Huang et al. (2019). It is inadequate at addressing real-world questions, where no context is readily provided for a system to find answers. Such open-ended questions require a system to identify answers in an enormous knowledge base (e.g., Wikipedia) that is computationally prohibitive to feed into question answering systems such as large language models (LLMs). Retrieval-augmented generative question answering (RAG-QA) becomes an effective tool to filter out massive amounts of noise and select only a few highly relevant passages for LLM-based QA models.
The wide applications of RAG-QA Gao et al. (2023) necessitate the evaluation of systems’ out-of-domain (OOD) performances, because a real-world system often confronts new data unseen during training. Existing popular benchmark datasets such as Natural Questions (NQ) Kwiatkowski et al. (2019) and TriviaQA Joshi et al. (2017) consist solely of Wikipedia or Web documents, which fall short at measuring OOD performances.
Robustqa Han et al. (2023) was the first dataset created to benchmark cross-domain robustness for RAG-QA. However, as illustrated by the yellow highlights in Figure 1, Robustqa follows NQ’s annotation format with short answer spans extracted from the documents. Such data format is not the most suitable reference answer to evaluate the current leading LLMs that typically generate long-form responses with multiple pieces of information combined in one coherent narrative Brown et al. (2020); Ouyang et al. (2022); OpenAI (2024). Consequently, token overlap metrics used in the extractive QA era Karpukhin et al. (2020); Izacard et al. (2024) will penalize unfairly on the additional supporting tokens generated by LLMs, resulting in extremely low overlap scores. As an example, Fig. 1’s extractive answers have poor Exact-Match or $F_{1}$ scores with the final long-form answer. To create a long reference answer, one could simply concatenate these short answers, but the synthesized answers are either incoherent or ill-formatted, as illustrated by examples in Sec. 3.1.
To address these drawbacks, we propose long-form RobustQA (Lfrqa) that integrates multiple short extractive answers into a coherent long-form answer. Figure 1 shows an annotation where three extractive answers are combined by annotators to create a comprehensive answer. Table 1 summarizes seven features in Lfrqa  that make it uniquely beneficial for RAG-QA evaluations. ASQA Stelmakh et al. (2022) and ELI5 Fan et al. (2019) are the most similar datasets to Lfrqa. However, they are either not directly annotated against the underlying corpus (thus, not RAG-QA), or rely on single-domain data, which is insufficient to benchmark systems’ cross-domain performances.
@@ 5 Experimental Setup
Table 4: Pairwise comparisons between Robustqa, Lfrqa, and LLMs. GPT-4: GPT-4-0125-preview; Mixtral: Mixtral-8x22b-Instruct; Llama-3: Llama-3-70b-Instruct. Answer generated based on top 5 passages using ColBERT-v2. MBE stands for the model-based evaluator. All numbers are % except for Pearson Corr. and Cohen’s Kappa. *** indicates strong correlation with p-values $\ll 0.001$.
@@ 6 Results and Analysis
Rag-qa Arena Ranking Rating 95% CI Votes Rating 95% CI Votes Rating 95% CI Votes
@@
6.1 LFRQA v.s. RobustQA
[…]
6.2 Quality of Model-based Evaluator
6.3 RAG-QA Arena
Dashboard leaders.  GPT-4o leads the dashboard, with GPT-4-turbo and Mixtral-8x22b-Instruct as close runners-up. GPT-4o performs the best for [BI], [LI], [RE] and [WR] domains, Mixtral-8x22b-Instruct leads in [FI] and [SC], and GPT-4-turbo champions in [TE].
These results raise a research question about the impact of prompt engineering. We emphasize that the goal of Rag-qa Arena is to propose a reliable evaluation framework, not to conduct extensive prompt engineering or model training to pursue the best RAG-QA system. We provide a dev set of Lfrqa, which can be leveraged in future research for model development purposes.
Elo rating.  Table 3 shows dashboard results of win and win+tie ratio against Lfrqa. We can further convert these pairwise comparisons into Elo ratings similar to Chatbot Arena Chiang et al. (2024). Table 5 reports our Elo ranking. The leftmost column shows Rag-qa Arena’s ranking based on win ratio: Lfrqa is ranked on the top followed by GPT-4o, GPT-4-turbo, etc. Column (A) uses the same data in Table 3, but here we only have comparisons between Lfrqa and other LLMs’ responses, i.e. there are no direct comparisons between LLMs’ responses. Thus, the total number of votes for are 176.7K for Lfrqa, and 16.1K for other LLMs. The new ranking based on Elo rating aligns with Rag-qa Arena, but the 95% confidence interval (CI) is not yet able to separate all different model pairs.
In Column (B), we add pairwise comparisons for all unique model pairs on 700 randomly sampled queries across seven domains (100 for each domain). This increases the votes for all LLMs compared to 23.1K. We continue to add 700 more queries in Column (C), which further increases the votes for all LLMs to 30.1K. We rank these pairs with the same LLM evaluator as in the main result table, and found that the ranking in (C) based on Elo rating aligns perfectly with Rag-qa Arena, and as we increase the pairs, the 95% CI can finally separate different models.
In general, the total added preference pairs are $\frac{N\times K\times(K-1)}{2}$, where $K$ is the number of models in Table 3, and $N$=700 and 1400 for Column (B) and (C), respectively. With 43.6% increase of total pairs (and thus the compute), the final ranking is identical with Rag-qa Arena based on win ratio, and only differs only slightly with the win+tie ranking in Table 3 by flipping the order of llama-3-70b and Command R+. These results present additional evidence that our approach of using Lfrqa only for pairwise comparisons is reliable. Furthermore, it reduces the computational costs from $\mathcal{O}(K^{2})$ to $\mathcal{O}(K)$ as we now only need to compare each LLM response once with the ground-truth in Lfrqa.
Impact of model sizes.  For the non-GPT LLM family, more parameters lead to better performances, but a larger increase in model sizes does not always indicate greater performance gains in our study. For example, the two Qwen1.5 models have the second-largest difference, but the lift from the smaller to the larger model is marginal. We leave more rigorous investigations to future research.
7 Discussion
7.1 LLM as Annotators
Citation
Completeness Accuracy Helpfulness
Lfrqa 90.8 88.9 48.1
GPT-4 75.3 65.5 35.2
Using large language models to provide annotations has been explored in previous works Tan et al. (2024). It could provide a more scalable solution than human annotations but can suffer from hallucination and accuracy issues that require human validations Huang et al. (2023). We also experimented with LLM as annotators before we start human annotations. We subsample 100 queries from Lfrqa and prompt GPT-4-0125-preview to follow the similar procedure in Sec. 3.1 to combine answers (Appendix Table 12). Then we request our data linguists to compare Lfrqa and GPT-4 annotations based on 1) Completeness: whether all Robustqa answers are integrated into the final answers; 2) Citation Accuracy: whether citations in answers pointing to the right documents; 3) Helpfulness: defined the same as in Sec. 4. Table 6 shows Lfrqa out-performs GPT-4 annotations by 15.5%, 23.4% and 12.9% for the three dimensions, respectively, suggesting human annotations are both valuable and necessary for our task.
7.2 Alternative Evaluation Approaches
Using retrieved passages.  Rag-qa Arena leverages only Lfrqa’s annotations as ground-truth to directly evaluate LLM responses, and we explain this design choice in Sec. 4 that Lfrqa consists of complete and coherent answers that can be viewed as high-quality summary of all available answers in the entire corpus. This enables us to not show retrieved passages as they 1) increase the input length and thus the latency of an evaluator; and 2) they could contain incorrect information due to retrieval error, which mislead evaluators.
Using Lfrqa as references.  We can also use Lfrqa’s annotations as references when constructing the prompt for pairwise evaluation. That is, we can potentially compare a pair of LLMs’ responses by comparing them both against the references in a single trial. However, this approach would still require the similar $\mathcal{O}(K^{2})$ pairs as in the Elo rating, which is not as efficient as our proposed Rag-qa Arena framework.
For these reasons, we do not adopt the above two evaluation approaches. It is conceivable that prompt engineering, in-context example selections and even task specific evaluator training could further enhance alignments with human judges. We leave them for future research efforts.
8 Related Work
Longform QA datasets  have been proposed in prior work. ELI5 Fan et al. (2019) and LongFact Wei et al. (2024) contain answers that are either not annotated directly on the corpus, and or not created by humans. Krishna et al. (2021) also points out that ELI5’s small validation set has significant leakage from its train set. ASQA Stelmakh et al. (2022) is the most similar data to our work, but its corpus is in the single Wikipedia domain. Lfrqa is by far the RAG-QA dataset with the most comprehensive long-form answers.
9 Conclusion
Limitations
Ethics Statement
Appendix A Appendix
@@
A.1 Dev Set of LFRQA
@@ A.3 Annotation Interface
A.4 Answer Generation Prompts
@@ A.5 Evaluation Interface
A.6 Human Evaluation Instructions
@@
A.7 Evaluation Model Selection
@@ A.8 No Answer Ratio
A.9 Pairwise Evaluation Prompts